3. A car moving with constant acceleration covers the distance between two points 60 m apart in 6 s. Its velocity as it passes the second point is 15 m/s.
a) What is its velocity at the first point?
b) What is the acceleration?
1
Expert's answer
2020-12-10T11:09:37-0500

a) We can find the velocity of the car at the first point from the kinematic equation:


"d=\dfrac{v_0+v}{2}t,""v_0=\dfrac{2d}{t}-v=\dfrac{2\cdot 60\ m}{6\ s}-15\ \dfrac{m}{s}=5\ \dfrac{m}{s}."

b) We can find the acceleration of the car from another kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{15\ \dfrac{m}{s}-5\ \dfrac{m}{s}}{6\ s}=1.66\ \dfrac{m}{s^2}."

Answer:

a) "v_0=5\ \dfrac{m}{s}."

b) "a=1.66\ \dfrac{m}{s^2}."